what is distance between two parallel lines given the equations
y=3x+2 and y=3x-8

Answer by Alan3354(21605)

what is distance between two parallel lines given the equations
y=3x+2 and y=3x-8
----------
Distance between y-intercepts = 10
Slope m = 3
d = 10*cos(arctan(m)) (I worked this out recently)
d = 3.16227766...
d = sqrt(10)